OpenStudy (anonymous):

Cherry has n cans of orange juice and 4 cans of grape juice. She choose 2 cans of juice at random and finds that the probability of getting two cans of the same juice is less than 1/2. Suggest two possible values of n. @ganeshie8

ganeshie8 (ganeshie8):

Cherry can get both orange juice or both grape juice : probability for getting both orange juice : \(\large \mathbb{\frac{n}{n+4} *\frac{n-1}{n-1+4} } = \frac{n}{n+4} *\frac{n-1}{n+3} \) probability for getting both grape juice : \(\large \mathbb{\frac{4}{n+4} *\frac{4-1}{n+4-1} } = \frac{4}{n+4} *\frac{3}{n+3} \) probability of getting same juice : \(\large \mathbb{\frac{n}{n+4} *\frac{n-1}{n+3} + \frac{4}{n+4} *\frac{3}{n+3}} \)

ganeshie8 (ganeshie8):

see if that makes sense so far

OpenStudy (anonymous):

yes...

ganeshie8 (ganeshie8):

since the probability is less than 1/2 : \(\large \mathbb{\frac{n}{n+4} *\frac{n-1}{n+3} + \frac{4}{n+4} *\frac{3}{n+3}} < \frac{1}{2}\) \(\large \mathbb{\frac{n(n-1) + 12}{(n+4)(n+3)} < \frac{1}{2} }\)

ganeshie8 (ganeshie8):

crossmultiplying and simplifying gives u : since the probability is less than 1/2 : \(\large \mathbb{n^2-9n + 12 < 0 }\)

ganeshie8 (ganeshie8):

simply pick any two values for n, that satisfy the above inequality

ganeshie8 (ganeshie8):

n = 2 , 3 may work ?

OpenStudy (anonymous):

wow, inequality....i hate this. But thanks! I will do it by myself now! :) if the answer is incorrect, i will message you. thanks a lot! the answer is "any integral values from 2 to 7"

ganeshie8 (ganeshie8):

n = 2,3,4,5,6,7 will work

OpenStudy (anonymous):

i see. thanks! After i finish this question i will go to sleep. it's 10:37 pm now. haha thanks a lot!

ganeshie8 (ganeshie8):

wow ! you're to the east ha... :3 have good sleep :)

OpenStudy (anonymous):

thanks. you too. I have done it! I got 1.6277<n<7.3723 therefore, 2<n<7 , and it's correct! thanks a lot haha